OpenStudy (anonymous):

There is a number that is 8 less than the sum of 20 + 12. Find the number. Let n = the number.

OpenStudy (mathstudent55):

Let n = number How do you express 8 less that something? it's something minus 8 8 less than the sum of 20 + 12 = 20 + 12 - 8 n = 20 + 12 - 8 Do the operations on the right side to get n.